OpenStudy (anonymous):

Convert the numeral to a numeral in base 10.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }= \frac{ 2 }{ 5 }*\frac{ 2 }{ 2 }=\frac{ 2 }{5*2 }\] hope that helps

OpenStudy (anonymous):

sorry it is 2*2 in the numerator

OpenStudy (anonymous):

55?

OpenStudy (anonymous):

From what I got sir

OpenStudy (anonymous):

\[\frac{ 4 }{ 10 }\]

OpenStudy (anonymous):

Okay 11

OpenStudy (anonymous):

\(\large 21_5=1*5^0+2*5^1=1+10=11_{10} \)

OpenStudy (anonymous):

11

OpenStudy (anonymous):

pretty sure that's right...

OpenStudy (anonymous):

That should be right

OpenStudy (anonymous):

it's rigth

OpenStudy (anonymous):

2 x (5^1)+1 x (5^0)= (2 x 5)+1=10+1= 11 in base 10